OpenStudy (anonymous):

find the next 4 terms of the sequence a1= -5 an= an-1-8

OpenStudy (anonymous):

a2 a3 a4 a5

OpenStudy (anonymous):

-13 19 27 -29 -37 -21 11

OpenStudy (anonymous):

match them up

OpenStudy (anonymous):

Given that \(a_1=-5\), we see by your sequence formula that \(a_2 = a_{2-1}-8 = a_1 - 8 = -5 - 8 = -13\) \(a_3 = a_{3-1}-8 = a_2 - 8 = -13 - 8 = -21\) \(a_4 = a_{4-1}-8 = a_3 - 8 =\ldots\) \(a_5 = a_{5-1}-8 = a_4 - 8 =\ldots\) Can you take things from here and find \(a_4\) and \(a_5\)? :-)

OpenStudy (anonymous):

not really its my first time doing this im confused could u walk me through it

OpenStudy (anonymous):

nvm i got it thanks :)

OpenStudy (anonymous):

Sure; the sequence provided to you is \(a_n = a_{n-1}-8\). We're told the value of this sequence when \(n=1\); in particular, that's the value \(a_1 = -5\). The next term in the sequence occurs when \(n=2\); so you want to plug \(n=2\) into your sequence formula to see that \(a_n = a_{n-1}-8 \implies a_2 = a_{2-1} - 8 \implies a_2 = a_1-8\). At this point, we note that we already know the value of \(a_1\), so we now substitute it into this equation to see that \(a_2 = a_1 - 8 = -5 - 8 = -13\). We repeat this same process for the third term; in this instance, \(n=3\). Thus, the sequence formula gives us \(a_n = a_{n-1}-8 \implies a_3 = a_{3-1}-8 \implies a_3 = a_2 - 8\). In the previous step though, we found that \(a_2 = -13\). We now substitute this into the sequence equation to see that \(a_3 = a_2 -8 = -13 - 8 = -21\). You now want to apply this process again to find \(a_4\) and \(a_5\). Does this clarify things?

OpenStudy (anonymous):

Oh, good to see you got it! XD